As an ambulance sounding a siren approaches a control point, the frequency of the siren is measured to be 360 Hz and as it passes and moves away, the frequency is presumed to be 320Hz.

 Explain why there is a difference and calculate the speed of the ambulance and the overall Doppler

shift that occurs


1
Expert's answer
2021-03-23T08:08:02-0400

Answer

Frequency can be written as when source is moving

"f_0=f_s(\frac{1}{1+\frac{v_{source}}{v_{sound}}}) \\320=360(\frac{1}{1+\frac{v_{source}}{343}})\\v_{source}=42.8m/s"